What is the shortest distance from the surface xy+12x+z^2=144 to the origin?

distance = ?
1
Expert's answer
2020-04-10T17:41:07-0400

Let P(x,y,z) be the point on the surface "xy + 12x + z^{2} = 144" . Then, distance between these two points is "d = \sqrt{x^2+y^2+z^2}". We have to find the point (x, y, z) which is at a minimum distance from origin. We now use Lagrange's multipliers to find the shortest distance.


Let "f(x,y,z) = d^{2} = x^{2} + y^{2} + z^{2}" and "g(x,y,z) = xy+12x+z^{2} = 144".

Let

"F(x,y,z) = f(x,y,z) + \lambda g(x,y,z)\\ = x^{2} + y^{2}+z^{2}+\lambda(xy+12x+z^{2}-144)"


To find the point of shortest distance we have to solve,

"\dfrac{\partial F}{\partial x} = 0; \dfrac{\partial F}{\partial y} = 0; \dfrac{\partial F}{\partial z} = 0~ and ~\dfrac{\partial F}{\partial \lambda} = 0."


"2x+\lambda(y+12) = 0\\
2y+\lambda x=0\\
2z+2\lambda z=0\\
xy+12x+z^2=144."


From the third equation, we get "z=0~~\text{or}~~ \lambda =-1". When "\lambda =-1", first and second equation becomes,

"2x-y=12\\
-x+2y=0."

Solving we get, "x= 8, y=4." From the constraint equation we get, "z = \pm 4."

The critical points are "(8,4,-4), (8,4,4)."


Now when z = 0, the constraint becomes, "xy+12x=144."

"x(y+12)=144\\
x=\dfrac{144}{y+12}."Using this in f(x,y), we get

"f(x,y) = (\dfrac{144}{y+12})^{2}+y^2."

Now we find the critical points of this one variable problem. We solve,

"\dfrac{\partial f}{\partial y} = 0\\
-\dfrac{2 \cdot 144^2}{(y+12)^3}+2y =0\\
y(y+12)^3-144^2=0\\
y^4+36y^3+432y^2+1728y-20736=0"


Solving this using online calculator, we get

"y=-9.366\pm10.973i\\
y=-21.83\\
y=4.5633"


Thus the critical points are, "(8.694,4.5633,0)" and "(-14.65,-21.83,0)" . The value of f at all critical points are,

"f(8,4,-4)=96\\
f(8,4,4)=96\\
f(8.694,4.5633,0)=96.41\\
f(-14.65,-21.83,0)=691.17"



Therefore, the shortest distance is "d = \sqrt{96} = 4\sqrt{6}" which is attained at "(8,4,\pm4)".